# REACT: A Real-Time Edge-AI Based V2X Framework for Accident Avoidance in Autonomous Driving System

### v1

@@ Abstract
Collisions caused by human error are the most common type of multi-vehicle crash, highlighting the critical need for autonomous driving (AD) systems to leverage cooperative perception through Vehicle-to-Everything (V2X) communication. This capability extends situational awareness beyond the limitations of onboard sensors. However, current transformer-based V2X frameworks suffer from limited generalization, shallow contextual reasoning, and reliance on mono-modal inputs. Vision-Language Models (VLMs) offer enhanced reasoning and multimodal integration but typically fall short of real-time performance requirements in safety-critical applications. This paper presents REACT, a real-time, V2X-integrated trajectory optimization framework built upon a fine-tuned lightweight VLM. REACT integrates a set of specialized modules that process multimodal inputs into optimized, risk-aware trajectories. To ensure real-time performance on edge devices, REACT incorporates edge adaptation strategies that reduce model complexity and accelerate inference. Evaluated on the DeepAccident benchmark, REACT achieves state-of-the-art performance, a 77% collision rate reduction, a 48.2% Video Panoptic Quality (VPQ), and a 0.57-second inference latency on the Jetson AGX Orin. Ablation studies validate the contribution of each input, module, and edge adaptation strategy.These results demonstrate the feasibility of lightweight VLMs for real-time edge-based cooperative planning and showcase the potential of language-guided contextual reasoning to improve safety and responsiveness in autonomous driving.
Keywords Traffic safety  $\cdot$ Edge AI  $\cdot$ V2X Communication  $\cdot$ Autonomous Driving  $\cdot$ Vision-Language Model
@@ 1 Introduction
Recent advances in Vision-Language Models (VLMs), extensions of Large Language Models (LLMs) with integrated visual processing capabilities [15], present promising solutions to overcome the limitations of conventional perception systems [16]. Modern traffic environments are dynamic, complex, and context-rich [17], demanding autonomous systems that can understand and respond to diverse multimodal inputs. VLMs address this challenge by combining semantic reasoning with visual context, enabling more informed and adaptive decision-making, outperforming conventional models [16]. First, VLMs excel in high-level contextual reasoning, interpreting spatiotemporal cues, anticipating agent intent, and understanding implicit traffic rules, which are essential for navigating scenarios like occluded merges, crowded intersections, or rare accident scenes [18, 19]. Second, VLMs inherently support multimodal integration, effectively fusing heterogeneous sensor data and textual inputs [20], which facilitates comprehensive perception and enables more accurate, adaptive, and timely decision-making in real-world driving scenarios [21, 22]. Third, VLMs demonstrate strong generalization through zero-shot learning [23, 24], reducing the dependency on large-scale labeled datasets and enhancing the robustness and reliability of perception systems. Finally, their ability to pair linguistic explanations with visual perception improves model interpretability [25, 26], supporting more efficient human-vehicle interactions by winning the trust of human drivers.
@@
  • Architectural Contribution: Introduction of REACT, a novel V2X-enabled framework that extends situational awareness beyond onboard sensing by leveraging cooperative communication to support safer and more informed autonomous decision-making.
  • Methodological Innovation: Design of a customized lightweight VLM that integrates multimodal sensor fusion, structured prompt optimization, and RTF to to enhance trajectory prediction accuracy, compatibility with existing planning systems, and applicability across diverse traffic conditions and planning tasks.
  • Edge Optimization: Implementation of prompt engineering and fine-tuning strategies tailored for edge-device deployment, ensuring real-time inference capabilities without sacrificing predictive performance.
@@ 3.2 System Configuration
[…]
@@ 3.3 Data Acquisition
As the first module of the REACT system shown in 2, the Data Acquisition stage collects both visual and non-visual inputs required for downstream reasoning. This module integrates raw sensor data and contextual information from vehicle and infrastructure sources into a unified pipeline.
@@ 3.5 Task Alignment
[…]
@@
[…]
@@ 3.6 Task Projection Enhancement
[…]
@@ 5.1 Experimental Setup
This study evaluates various edge adaptation strategies, as detailed in the ablation studies. Based on the experimental outcomes, the optimal configuration utilizes 16-bit quantization, a token length of 1500, and a BEV resolution of $\times 64$ 64. Subsequent experiments adopt this optimized configuration to ensure optimal performance.
@@ 5.5.2 Effect of Edge Adaption Strategies
This section investigates how various edge adaptation strategies on AGX affect the runtime efficiency and output quality of the REACT framework. The Table 4 reveals that token length reduction and BEV size compression are the most effective in reducing inference latency with minimal VPQ degradation. While 8-bit quantization generally offers lower latency, certain configurations favor 16-bit due to hardware-specific optimizations. SDPA consistently improves speed but shows diminishing returns under lightweight settings. Based on the trade-off between runtime and predictive quality, the optimal configuration includes 16-bit quantization, a 1500-token input length, a $\times 64$ 64 BEV size, and SDPA. The following paragraphs provide a detailed breakdown of how each individual strategy contributes to this balance.

### v2

# Edge-Based Multimodal Sensor Data Fusion with Vision Language Models (VLMs) for Real-time Autonomous Vehicle Accident Avoidance

@@ Abstract
Autonomous driving (AD) systems relying solely on onboard sensors may fail to detect distant or obstacle hazards, potentially causing preventable collisions; however, existing transformer-based Vehicle-to-Everything (V2X) approaches, which mitigate AD sensing limitations, either lack effective multimodal fusion and reasoning or struggle to meet real-time performance requirements under complex, high-dimensional traffic conditions. This paper proposes the Real-time Edge-based Autonomous Co-pilot Trajectory planner (REACT), a V2X-integrated trajectory optimization framework for AD based on a fine-tuned lightweight Vision-Language Model (VLM). REACT integrates infrastructure-provided hazard alerts with onboard sensor data, capturing intricate surrounding traffic dynamics and vehicle intents through visual embeddings, interpreting precise numerical data from symbolic inputs, and employing contextual reasoning to generate optimized, safety-oriented trajectories. To ensure robust real-time deployment on edge devices, REACT innovatively employs Residual Trajectory Fusion (RTF) design and specialized edge-adaptation strategies to reduce model complexity and improve inference efficiency. Evaluated on the DeepAccident benchmark, REACT achieves state-of-the-art performance, a 77% collision rate reduction, a 48.2% Video Panoptic Quality (VPQ), and a 0.57-second inference latency on the Jetson AGX Orin. Ablation studies validate the contribution of each input, module, and edge adaptation strategy. These results highlight the effectiveness of lightweight VLMs in enabling real-time cooperative planning on edge platforms and underscore the potential of language-guided contextual reasoning for improving traffic safety and responsiveness.
Keywords Large Language Model  $\cdot$ Multi-modal Data Fusion  $\cdot$ Spatial-temporal Alignment  $\cdot$ Transportation Safety  $\cdot$ Connected and Autonomous Vehicles  $\cdot$ Edge Computing
@@ 1 Introduction
Recent advances in Vision-Language Models (VLMs), extensions of Large Language Models (LLMs) with integrated visual processing capabilities [15], present promising solutions to overcome the limitations of conventional perception systems [16, 17]. Modern traffic environments are dynamic, complex, and context-rich [18], demanding autonomous systems that can understand and respond to diverse multimodal inputs. VLMs address this challenge by combining semantic reasoning with visual context, enabling more informed and adaptive decision-making, outperforming conventional models [16]. First, VLMs excel in high-level contextual reasoning, interpreting spatiotemporal cues, anticipating agent intent, and understanding implicit traffic rules, which are essential for navigating scenarios like occluded merges, crowded intersections, or rare accident scenes [19, 20]. Second, VLMs inherently support multimodal integration, effectively fusing heterogeneous sensor data and textual inputs [21], which facilitates comprehensive perception and enables more accurate, adaptive, and timely decision-making in real-world driving scenarios [22, 23]. Third, VLMs demonstrate strong generalization through zero-shot learning [24, 25], reducing the dependency on large-scale labeled datasets and enhancing the robustness and reliability of perception systems. Finally, their ability to pair linguistic explanations with visual perception improves model interpretability [26, 27], supporting more efficient human-vehicle interactions by winning the trust of human drivers.
@@
  • Architectural Contribution: This paper introduces REACT, a V2X-integrated autonomous trajectory planning framework built upon a customized VLM to enable safer and more context-aware autonomous driving.
  • Methodological Innovation: This study designs a lightweight VLM-based sensor fusion system that integrates multimodal information via spatial-temporal alignment and spatial grounding enhancement, with only 512M parameters. Furthermore, an innovative structure Residual Trajectory Fusion (RTF) is proposed to improve both system computation efficiency and accuracy for collision risks mitigation.
  • Edge Optimization: REACT incorporates targeted edge adaptation strategies that streamline model complexity, enabling real-time deployment on resource-constrained edge devices without compromising predictive quality.
  • Practical Applicability: The proposed system is validated under diverse traffic scenarios, environmental conditions, and infrastructure configurations, demonstrating robustness and adaptability in real-world deployment settings.
@@ 3.2 System Configuration
To provide a unified spatial reference, all vehicle states are defined within a common coordinate system centered at the roadside unit (RSU), denoted as $F_{R}=\{\hat{\mathbf{x}}_{\text{up}},\,\hat{\mathbf{y}}_{\text{right}},\,\hat{\mathbf{z}}_{\text{up}}\}$. In this frame, each vehicle $i\in\mathcal{I}$ in vehicle set $\mathcal{I}$ at discrete timestep $k\in\mathcal{K}$ in time range $\mathcal{K}$ is characterized by Eq. 1, where $(x,y,z)$ represents the longitudinal, lateral, and vertical positions relative to the RSU origin, $(v_{x},v_{y})$ denotes velocities in the RSU reference plane, and $\psi$ is the yaw angle representing the vehicle’s heading relative to the longitudinal axis.
@@ 3.3 Data Acquisition
As the first module of the REACT system shown in Figure 2, the Data Acquisition stage collects both visual and non-visual inputs required for downstream reasoning. This module integrates raw sensor data and contextual information from vehicle and infrastructure sources into a unified pipeline.
@@ 3.5 Task Alignment
To furnish the VLM with explicit spatial grounding, real-world coordinates are transferred and overlaid onto the BEV feature map $\mathbf{B}_{t}\in\mathbb{R}^{C\times H\times W}$, which is defined on a discrete image grid $(u,v)\in\mathcal{G}\subset\{0,\ldots,W{-}1\}\times\{0,\ldots,H{-}1\}$, where $(u,v)$ denotes a pixel in the BEV map and $(u_{0},v_{0})$ is the ego-vehicle pixel (the anchor). Let $r$ be the metre-per-pixel scale. The ego-centred coordinates $(x,y$ corresponding to any BEV pixel are obtained by Eq. 14.
@@
To ensure consistent spatial grounding for downstream VLM reasoning, all context elements in $\mathcal{C}_{t}$ are represented in the ego–centered frame. Since each BEV is already ego–referenced, the hazard set $\mathcal{H}_{t}$, the effective navigation waypoints $\mathcal{N}_{\text{eff}}$, and the ego trajectory history $\mathcal{S}_{\text{ego}}$ are translated into the same frame by Eq. 16, where $(x,y)$ denotes any global position from $\mathcal{H}_{t}$, $\mathcal{N}_{\text{eff}}$, or $\mathcal{S}_{\text{ego}}$, and $(x_{e,k},y_{e,k})$ is the current ego position.
@@ 3.6 Task Projection Enhancement
For each training instance, the training set is prepared that includes: (1) a predefined instruction $\mathcal{P}_{Inst}$ explaining the optimization goal, (2) a multimodal prompt $P=\{P_{t},P_{v}\}$ representing the scenario context, (3) a human-written CoT-style reasoning trace $\mathcal{R}$ that describes the step-by-step planning rationale, and (4) a human-labeled answer. Together, they form the input set as Eq. 20, where $\|$ denotes concatenation across instruction, symbolic, visual, and CoT components.
@@ 5.1 Experimental Setup
This study evaluates various edge adaptation strategies, as detailed in the ablation studies. Based on the experimental outcomes, the optimal configuration utilizes 16-bit quantization, a token length of 1500, and a BEV resolution of $64\times 64$. Subsequent experiments adopt this optimized configuration to ensure optimal performance.
@@ 5.5.2 Effect of Edge Adaption Strategies
This section investigates how various edge adaptation strategies on AGX affect the runtime efficiency and output quality of the REACT framework. The Table 4 reveals that token length reduction and BEV size compression are the most effective in reducing inference latency with minimal VPQ degradation. While 8-bit quantization generally offers lower latency, certain configurations favor 16-bit due to hardware-specific optimizations. SDPA consistently improves speed but shows diminishing returns under lightweight settings. Based on the trade-off between runtime and predictive quality, the optimal configuration includes 16-bit quantization, a 1500-token input length, a $64\times 64$ BEV size, and SDPA. The following paragraphs provide a detailed breakdown of how each individual strategy contributes to this balance.
@@ 6 Conclusion and Discussion
7 Funding
The authors disclosed no financial support for the research, authorship, and/or publication of this article.
8 Declaration of Generative AI and AI-assisted Technologies in the Writing Process
During the preparation of this work the author(s) used OpenAI’s GPT-4o language model solely in order to polish grammar and improve the clarity of writing during manuscript preparation. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
9 Author Contributions
Fengze Yang led the conceptualization, methodology design, software development, and original draft writing. Bo Yu conducted the experiments and prepared the baseline models based on the dataset. Zhou Yang contributed to idea refinement, paper structuring, and writing. Zhengzhong Tu contributed to model fine-tuning, data processing, and manuscript writing. Xuewen Luo contributed to the literature review, data processing, and manuscript polishing. Chenxi (Dylan) Liu (corresponding author) oversaw project administration and supervision, and contributed to manuscript review and editing.
10 Declaration of Conflicting Interests
All authors declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.